Grade 12Inorganic Chemistry

Henry's law constant of oxygen is 1.4 × 10^-3 mol.lit^-1.atm^-1 at 298 K .How much of oxygenis dissolved in 100 ml at 298 K when partial pressure of oxygen is 0.5 atm

Sanketbhole
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
KH = C/P
1.4 x 10-3 = C/0.5

Therefore C = 1.4 x 0.5 x 10-3 

No. of moles of O2 in 100 mL
=  (1.4 x 0.5 x 10-3) 100/1000
= 1.4 x 0.5 x 10-4

Weight of O2 in 100 mL
=  1.4 x 0.5 x 10-4  x 32 g
= 2.24 x 10-3 g
= 2.24 mg